OpenStudy (sam_ovo):

radical form of 11w^-2.5

OpenStudy (sam_ovo):

\[11w ^{-2.5}\]

OpenStudy (unklerhaukus):

\[\large a^{n/m}=\sqrt[m]{a^n}\]

OpenStudy (unklerhaukus):

can you express the index (the -2.5) as a fraction?

OpenStudy (sam_ovo):

-5/2

OpenStudy (unklerhaukus):

good,

OpenStudy (sam_ovo):

\[11/\sqrt[2]{w^5}\]

OpenStudy (unklerhaukus):

yeah you've got rid of the minus too good work,

OpenStudy (sam_ovo):

thank you!

OpenStudy (unklerhaukus):

final step is remove the little 2 \[\sqrt[2] x=\sqrt x\](the 2 is implied)

OpenStudy (sam_ovo):

right lol